# Cosmic Ray Diffusion in the Turbulent Interstellar Medium: Effects of Mirror Diffusion and Pitch Angle Scattering

### v1

@@ Abstract
Cosmic rays (CRs) interact with turbulent magnetic fields in the intestellar medium, generating nonthermal emission. After many decades of studies, the theoretical understanding of their diffusion in the ISM continues to pose a challenge. This study numerically explores a recent prediction termed ”mirror diffusion” and its synergy with traditional diffusion mechanism based on gyroresonant scattering. Our study combines 3D MHD simulations of star-forming regions with test particle simulations to analyze CR diffusion. We demonstrate the significance of mirror diffusion in CR diffusion parallel to the magnetic field, when the mirroring condition is satisfied. Our results support the theoretical expectation that the resulting particle propagation arising from mirror diffusion in combination with much faster diffusion induced by gyroresonant scattering resembles a Levy-flight-like propagation. Our study highlights the necessity to reevaluate the diffusion coefficients traditionally adopeted in the ISM based on gyroresonant scattering alone. For instance, our simulations imply a diffusion coefficient $\sim 10^{27}cm^{2}/s$ for particles with a few hundred TeV within regions spanning a few parsecs around the source. This estimate is in agreement with gamma-ray observations, which shows the relevance of our results for understanding of diffuse gamma-ray emission in star-forming regions.
@@ 1 Introduction
Cosmic rays (CRs) undergo the interactions with magnetic fluctuations and change the direction of their trajectories as they diffuse through the magnetized interstellar medium (ISM). The amount of time these particles spend during their journey through the ISM plays a crucial role in elucidating phenomena such as propagation, acceleration, and non-thermal emission. The commonly referred approach to studying CR diffusion is the quasi-linear theory (QLT Jokipii, 1966; Schlickeiser, 2002), which evaluates the diffusion of CR considering small perturbations in the average magnetic field.
Accounting for the actual properties of MHD turbulence (see Beresnyak & Lazarian, 2019) is essential for understanding the pitch-angle scattering. For instance, Yan & Lazarian (2002, 2004) demonstrated that fast modes are more important compared to the other modes, i.e., Alfvén and slow modes, for gyroresonant scattering of CRs by turbulence injected at the scale height of the spiral galaxies. This happens in the ionized medium in most of galactic environments, even though the fast modes are subject to collisionless damping. The variations of magnetic field strength, arising, for instance, from the large scale slow modes, induce the modification of QLT and change the scattering rate (Yan & Lazarian, 2008; Lynn et al., 2012; Xu & Lazarian, 2018).
@@
A long-standing theoretical difficulty with scattering is the $90^{\circ}$ problem (Jokipii, 1966). QLT falls short in accounting for the scattering when the particle’s pitch angle approaches $90^{\circ}$. Earlier investigations have explored phenomenological models of turbulence in conjunction with QLT (see e.g. Matthaeus et al., 1990; Giacalone & Jokipii, 1999; Shalchi, 2016). However, these models have faced challenges from observations and simulations (Qin et al., 2002; Gabici et al., 2019). Inevitably, the solutions ought to rely on thoroughly validated models of MHD turbulence. Such models can significantly alter pitch angle scattering, acceleration, and spatial diffusion of CRs compared to previous ad hoc approaches (e.g. Yan & Lazarian, 2002, 2004; Brunetti & Lazarian, 2007; Beresnyak et al., 2011; Xu & Yan, 2013; Cohet & Marcowith, 2016; Medina-Torrejón et al., 2021; Kadowaki et al., 2021; Fornieri et al., 2021; Kempski et al., 2023; Lemoine, 2023; Medina-Torrejón et al., 2023; Gao & Zhang, 2024).
@@
The super-diffusive behavior of turbulent magnetic field lines (Lazarian & Vishniac, 1999; Xu & Yan, 2013; Lazarian & Yan, 2014; Beresnyak, 2013) and particle interaction with magnetic mirrors (Kulsrud & Pearce, 1969; Xu & Lazarian, 2020) have been combined in Lazarian & Xu (2021) (henceforth LX21) to predict a new diffusion mechanism which was termed “mirror diffusion”. LX21 demonstrated that in the presence of superdiffusion of magnetic field lines, the probability of CRs getting trapped and bouncing back and forth between the same magnetic mirrors is negligibly small. CRs do not retrace their trajectories and bounce every time from different mirrors. As the mirroring causes particles to reverse their moving directions, the mirror diffusion naturally solves the 90 degree problem and is a much slower diffusion process compared to the diffusion induced by gyroresonant scattering alone.
@@
Naturally, this new process affects only the particles with Larmor radius smaller than the size of the magnetic mirrors and a pitch angle larger than the critical angle $\alpha_{pitch,c}$. This is equivalent to say that $\mu=\cos\alpha_{pitch}$ is less than some critical value which depends on the amplitude of magnetic compressions and the efficiency of pitch-angle scattering (LX21).
@@
The magnetic compressions induced by the cascade of fast and slow modes form a hierarchy of amplitudes from large compressions $\delta B_{L}$ at the injection scale $L$ to the smallest ones $\delta B_{d}$ at the dissipation scale $l_{d}$. These magnetic compressions induce the hierarchy of magnetic mirrors of different sizes. CRs with the largest $\mu=\mu_{c}$ bounce from large-scale magnetic compressions, while CRs with smaller $\mu$ bounce from magnetic compressions at smaller scales. As mirror diffusion preserves the first adiabatic invariant $(1-mv_{\bot}^{2})/B$ (Kulsrud & Pearce, 1969), the initial value of $\mu$ does not changed, and thus, the diffusion coefficient for mirror diffusion alone is a function of $\mu$. Effective mirroring requires that the CR gyroradius is smaller than the perpendicular extension of the compressible magnetic fluctuation. As $\mu_{c}$ where the transition between mirror diffusion and scattering diffusion takes place depends on CR energy (LX21), thus, the overall CR diffusion depends on its Larmor radius, i.e., on its energy.
@@
In highly compressible environments like molecular clouds, we expect fast modes to dominate the mirroring effect (Xu & Lazarian, 2020; Xu, 2021). The resulting diffusion coefficient for mirror diffusion shows a strong $\mu$-dependence (LX21 and Zhang & Xu, 2023). The dependence on $\mu$, makes the mirror diffusion very different from the classical resonance scattering diffusion that isotropizes $\mu$. If we are interested in the total CR diffusion flux, we need information on the distribution of $\mu$, a new element for the CR propagation paradigm. It is reasonable to assume that CRs have the isotropic $\mu$-distribution near the source. With this assumption, considering mirror diffusion, Xu (2021) estimated the spectrum of CRs in molecular clouds (MCs) with highly compressible turbulence subject to interactions with supernova remnants (SNR) and found that it leads to a steeper high-energy gamma-ray spectrum which can naturally explain observed systems like IC 443 and W44.
@@
Certain analytical predictions from LX21 were also validated successfully in Zhang & Xu (2023). However, the crucial importance of magnetic bouncing and mirror diffusion motivates us for more detailed studies of the processes. In this work, we investigate the mirror diffusion of charged particles in a highly compressible environment using 3D-MHD simulations of an MC-like environment as the background and test particle propagation models to evaluate the diffusion behavior of the particles aiming at testing numerically the aforementioned predictions of Xu & Lazarian (2020) and LX21.
@@ 2.1 Background MHD turbulence
Figure 1 shows the energy spectrum of the turbulent magnetic field for the background MHD model, after turbulence has reached a nearly steady state (see Barreto-Mota et al., 2021, for details). The blue line shows the results for the complete background magnetic field while the orange one shows the same model with a smoothed background magnetic field, where the smaller scale perturbations have been suppressed. To smooth the small scale perturbations, we calculated the 3D Fourier transform of each component of the magnetic field, set to zero the values corresponding to $k>5$, and then transformed it back to physical space. The green dashed line shows the Kolmogorov scaling with $k^{-5/3}$ for comparison. We see that a substantial part of the inertial range of the magnetic field in this smoothed model is eliminated. We will employ both models for the CR propagation analysis. The smoothing of the magnetic field reduces the effects of scattering on the propagation of particles, allowing us to focus on the effects of mirroring diffusion.
Figure 1: Magnetic energy spectrum of the magnetic field components used in our models. The blue line shows the original complete magnetic energy spectrum (Complete models as referred in the text). The orange line shows that for the smoothed magnetic field (see text for details).
@@ 2.2 Test Particle Simulations
To study the CR diffusion in the turbulent magnetic field generated in the aforementioned model, we employ the test particle code developed by Beresnyak et al. (2011) (see also Xu & Yan, 2013; Hu et al., 2022). The particles are allowed to diffuse through the nearly steady state background turbulent magnetic field of the simulated model described in Section 2.1.
@@
where $R_{L}$ is its Larmor radius, $m$ its mass, $q$ its charge, and $\Omega$ its gyrofrequence.In this study, we consider the particle velocity approximately given by the light speed $c$. We take particles with constant Larmor radius and initial $v_{\perp,0}\sim c$.
The code uses a Burlisch-Stoer method to trace the trajectories of test particles and uses an adaptive time-step. At each time step, the magnetic field at the grid points is interpolated to the test particles position using a cubic spline routine (Press et al., 1986; Xu & Yan, 2013).
Given that the MHD turbulent models are scale free and the energy of the particles depend on the properties of the simulated box (see Section 4), one can scale the system knowing the size of the domain $L_{0}$, the mean density and the sound speed as well as the scale independent parameters $M_{A}$ and $M_{s}$ that are defined at the injection scale of turbulence.
@@ 2.3 Initial setup
Each model is characterized by two free parameters, namely, the initial pitch angle ($\theta_{0}$) between the particles velocity and the local magnetic field, which we consider to be either $\mu_{0}=\cos\theta_{0}=0.2$ or $0.8$, and the Larmor radius ($R_{L}$), $R_{L}/L_{0}=0.03,0.06$ and $0.10$, in terms of the number of cells of the background simulation, $R_{L}(\#cells)=15,30$, and $50$, respectively. We note that particles are not accelerated and their energy stays the same in our simulations, we focus only in their diffusion. Each model name is labeled with these two parameters in Table 1. As described in section 2.1, we also consider two sets for the magnetic field configuration. One with the entire background turbulent magnetic field distribution, and another where the small-scale perturbations of the field were suppressed for $k_{cut}=5$, which corresponds to a dissipative scale of about $\sim 0.2L_{0}$ ($\sim 100$ grid cells) [^1] Values of $k_{cut}>5$ were tested, but the results showed little to no notable differences with regard to the model with the complete distribution of the magnetic field.. These latter models are labeled $k5$ in Table 1. The dissipation scale of the complete distribution of magnetic field is around $k\sim 250$, which corresponds to $0.004L_{0}$ approximately ($\lesssim 3$ grid cells). Models with the first background are initialized with 1000 particles, while the ones with the second background are initialized with 3000 particles. This difference will be discussed in the next sections.
@@ 3.1 Particle trajectory and mirroring identification
Figure 2: Example of a test particle trajectory (represented by solid colored circles) with $R_{L}/L_{0}\simeq 0.03$ and $\mu_{0}=0.2$ in the background with complete distribution of magnetic field. Initially, the particle travels upwards parallel to the magnetic field (black circles), reentering the domain from the bottom face until it encounters magnetic mirrors that trap it for a while (dark purple to orange circles), diffusing perpendicularly to the field lines until it escapes again traveling downwards and reentering through the upper region of the domain, finding another mirror at the end of its trajectory (bright yellow circles). A few field lines are shown in red, white, and blue for visual distinction only. The color scales on the right hand side show the time elapsed in the particle trajectory between 850 and 1450 (same units as in Figure 3), and the density of the background (in code units), respectively. See Figure 3 and text for further details. We have plotted only a section of the cubic box in order to make the particle trajectory clearer.
@@
Figure 2 depicts an example of a particle trajectory in the original MHD turbulence simulation. A characteristic behavior observed in every model is long flights interrupted by successive bounces, where the particle interacts with different magnetic mirrors (Levy-flght-like propagation). As discussed in LX21, only particles with pitch angles above a critical value should interact with magnetic mirrors present along their path (Equation 4). When these conditions are met, the particle reverses its moving direction and bounces between mirrors. In the meantime, the particle also undergoes superdiffusion in the direction perpendicular to the mean magnetic field as a result of the turbulent reconnection of the magnetic fields. Because of the perpendicular superdiffusion, the particle does not remain trapped by mirrors, but diffusively move along the magnetic field when interacting with different mirrors. When the pitch angle becomes sufficiently small due to the pitch-angle scattering, the particle can escape from mirrors and undergo the scattering diffusion. With the stochastic change of pitch angle due to the gyroresonant scattering, the particle stochastically experiences both the mirror diffusion and scattering diffusion.
@@
Figure 3: The first panel depicts the time evolution of the cosine of the pitch angle ($\mu$) for the same particle shown in Figure 2, for complete background magnetic field distribution. The gradient color scheme depicts the portion of the particle trajectory as in Figure 2. Gray line shows the complete variation of $\mu$ over time and the blue line shows the same, but without the fastest variations of $\mu$. The middle panel shows the displacement of the particle parallel to the magnetic field with regard to its initial position, and the bottom panel shows $\zeta/\zeta_{0}$ (Equation 3) over time. Red circles stand for the identified bounces. Black triangles are points that are not considered as bounces due to mirror diffusion (see text for details). Time is normalized by the initial value of the girofrequency $\Omega=c/R_{L}$ (Equation 2), being $c=1$ in the code units.
@@
In Figure 3, we show the time evolution of $\mu$ (top panel), the distance traveled parallel to the magnetic field versus time (middle panel), and $\zeta/\zeta_{0}$ versus time, where $\zeta$ is the adiabatic invariant of the particle, which is conserved when the particle is bouncing in a mirror:
@@
where $\zeta_{0}=\zeta(t=0)$ for the same particle shown in Figure 2.
As an example, Figure 2 shows just a part of the particle’s trajectory. The respective time interval is shown in Figure 3 with the same color coding. Considering this time interval, the second panel of Figure 3 shows that the particle travels upward until it interacts with a mirror, staying more or less in the same height until it is able to continue its trajectory downwards after several consecutive mirroring events. In Figure 2 we see that during these several consecutive bounces, the particle continues to move perpendicular to the magnetic field, interacting with different mirrors. In the lower panel of Figure 3, the respective time interval also shows a nearly constant adiabatic invariant, which reinforces the fact that these are mirroring events.
The peaks right before each bounce point (i.e., when $\mu=0$) in Figure 3, are identified as the beginning of a bounce $\mu_{b}$ is the value of $\mu$ for each of these peaks. $\mu_{b}$ reflects the angle with which the particle starts to interact with a magnetic mirror. In other words, if $\mu_{b}>\mu_{c}$ the particle should not interact with the mirror, where $\mu_{c}=\cos{\alpha_{pitch,c}}$ is the critical value of the pitch angle cossine below which particle mirror diffusion occurs, as defined in Section 1.
@@
To identify the peaks in $\mu$, we first filter the values of $\mu$ along time using a Butterworth digital filter (Virtanen et al., 2020), which helps us to follow the average behavior of $\mu$ along time. Note that there is a fast variation of $\mu$and that not all bounce points correspond to $\zeta/\zeta_{0}$ constant. This variation is reduced comparing the first panels of Figures 3 and 4. Figure 4 shows the information presented in Figure 3, but for a particle propagating in a smoothed background distribution of the magnetic field. In Figure 3, top, the gray line (original data) has a component that partially hides the average behavior of the particle which is depicted by the blue line (filtered data). In Figure 4, since the magnetic field background is smoothed and scattering is already reduced, the peaks are more easily identifiable.

### v3

@@ Abstract
Cosmic rays (CRs) interact with turbulent magnetic fields in the intestellar medium, generating nonthermal emission. After many decades of studies, the theoretical understanding of their diffusion in the ISM continues to pose a challenge. This study numerically explores a recent prediction termed “mirror diffusion” and its synergy with traditional diffusion mechanism based on gyroresonant scattering. Our study combines 3D MHD simulations of star-forming regions with test particle simulations to analyze CR diffusion. We demonstrate the significance of mirror diffusion in CR diffusion parallel to the magnetic field, when the mirroring condition is satisfied. Our results support the theoretical expectation that the resulting particle propagation arising from mirror diffusion in combination with much faster diffusion induced by gyroresonant scattering resembles a Levy-flight-like propagation. Our study highlights the necessity to reevaluate the diffusion coefficients traditionally adopted in the ISM based on gyroresonant scattering alone. For instance, our simulations imply a diffusion coefficient $\sim 10^{27}cm^{2}/s$ for particles with a few hundred TeV within regions spanning a few parsecs around the source. This estimate is in agreement with gamma-ray observations, which shows the relevance of our results for understanding of gamma-ray emission in star-forming regions.
@@ 1 Introduction
Cosmic rays (CRs) undergo the interactions with magnetic fluctuations and change the direction of their trajectories as they diffuse through the magnetized interstellar medium (ISM). The amount of time these particles spend during their journey through the ISM plays a crucial role in elucidating phenomena such as propagation, acceleration, and non-thermal emission. The quasi-linear theory (QLT) (Jokipii, 1966; Schlickeiser, 2002), as a commonly used approach for studying CR diffusion, analyzes diffusion by considering small perturbations superimposed on the mean magnetic field.
Accounting for the actual properties of MHD turbulence (see Beresnyak & Lazarian, 2019) is essential for understanding the pitch-angle scattering. For instance, Yan & Lazarian (2002, 2004) demonstrated that fast modes are more important compared to the other modes, i.e., Alfvén and slow modes, for gyroresonant scattering of CRs by turbulence injected at the scale height of the spiral galaxies. This happens in the ionized medium in most of galactic environments, even though the fast modes are subject to collisionless damping. The variations of magnetic field strength, arising, for instance, from the turbulent compressions, induce the modification of QLT and change the scattering rate (Yan & Lazarian, 2008; Lynn et al., 2012; Xu & Lazarian, 2018).
@@
A long-standing theoretical difficulty with resonant scattering is the $90^{\circ}$-problem (Jokipii, 1966). QLT falls short in accounting for the scattering when the particle’s pitch angle approaches $90^{\circ}$. Resonant scattering alone fails in reversing the moving direction of particles. Earlier investigations have explored various phenomenological models of turbulence in conjunction with QLT (see e.g. Matthaeus et al., 1990; Giacalone & Jokipii, 1999; Shalchi, 2016). However, these models have faced challenges from observations and simulations (Qin et al., 2002; Gabici et al., 2019). Inevitably, the solutions ought to rely on thoroughly validated models of MHD turbulence. Such models can significantly alter pitch angle scattering, acceleration, and spatial diffusion of CRs compared to previous ad hoc approaches (e.g. Yan & Lazarian, 2002, 2004; Brunetti & Lazarian, 2007; Beresnyak et al., 2011; Xu & Yan, 2013; Cohet & Marcowith, 2016; Medina-Torrejón et al., 2021; Kadowaki et al., 2021; Fornieri et al., 2021; Kempski et al., 2023; Lemoine, 2023; Medina-Torrejón et al., 2023; Gao & Zhang, 2024).
@@
The super-diffusive behavior of turbulent magnetic field lines (Lazarian & Vishniac, 1999; Beresnyak, 2013; Eyink et al., 2013; Xu & Yan, 2013; Lazarian & Yan, 2014; Hu et al., 2022; Maiti et al., 2022; Zhang & Xu, 2023) and particle interaction with magnetic mirrors (Kulsrud & Pearce, 1969; Xu & Lazarian, 2020) have been combined in Lazarian & Xu (2021) (henceforth LX21) to predict a new non-resonant diffusion mechanism which was termed “mirror diffusion”. LX21 demonstrated that in the presence of superdiffusion of magnetic field lines, the probability of CRs getting trapped and bouncing back and forth between the same magnetic mirrors is negligibly small. CRs do not retrace their trajectories and bounce every time from different mirrors. As the mirroring causes particles to reverse their moving directions, the mirror diffusion naturally solves the 90 degree problem and is a much slower diffusion process compared to the diffusion induced by gyroresonant scattering alone.
@@
Naturally, this new process affects only the particles with Larmor radius smaller than the size of the magnetic mirrors and a pitch angle larger than the critical angle $\theta_{pitch,c}$. This is equivalent to say that $\mu=\cos\theta_{pitch}$ is less than some critical value at the balance between mirroring and scattering which depends on the amplitude of magnetic compressions and the efficiency of pitch-angle scattering (LX21).
@@
The magnetic compressions induced by the cascade of fast and slow modes form a hierarchy of amplitudes from large compressions $\delta B_{L}$ at the injection scale $L$ to the smallest ones $\delta B_{d}$ at the dissipation scale $l_{d}$. These magnetic compressions induce the hierarchy of magnetic mirrors of different sizes. CRs with the largest $\mu=\mu_{c}$ bounce from large-scale magnetic compressions, while CRs with smaller $\mu$ bounce from magnetic compressions at smaller scales. As mirror diffusion preserves the first adiabatic invariant $\gamma mv^{2}(1-\mu^{2})/2|\textbf{B}|$ (Kulsrud & Pearce, 1969), the initial value of $\mu$ does not stochastically change, but changes between ±$\pm\mu$, and thus, the diffusion coefficient for mirror diffusion alone is a function of $\mu$. Effective mirroring requires that the CR gyroradius is smaller than the parallel extension of the compressible magnetic fluctuation. As $\mu_{c}$ where the transition between mirror diffusion and scattering diffusion takes place depends on CR energy (LX21), thus, the overall CR diffusion depends on its Larmor radius, i.e., on its energy, but with the energy dependence different from the scattering diffusion alone [^1] Note that mirror diffusion and scattering diffusion are two separate mechanisms. Calling it “mirror scattering” would be conceptually incorrect..
@@
In highly compressible environments like molecular clouds, we expect fast modes to dominate the mirroring effect (Xu & Lazarian, 2020; Xu, 2021). The resulting diffusion coefficient for mirror diffusion shows a strong $\mu$-dependence (LX21 and Zhang & Xu, 2023). The dependence on $\mu$, makes the mirror diffusion very different from the classical resonance scattering diffusion that isotropizes $\mu$. If we are interested in the total CR diffusion flux, we need information on the distribution of $\mu$, a new element for the CR propagation paradigm. It is reasonable to assume that CRs have the isotropic $\mu$-distribution near the source. With this assumption, considering mirror diffusion, Xu (2021) estimated the spectrum of CRs in molecular clouds (MCs) with highly compressible turbulence subject to collisions with supernova remnants (SNR) and found that it leads to a steeper high-energy gamma-ray spectrum which can naturally explain observed systems like IC 443 and W44.
@@
Certain analytical predictions from LX21 were also validated successfully in Zhang & Xu (2023). However, the crucial importance of magnetic bouncing/mirroring and the resulting mirror diffusion motivates us for more detailed studies of the processes. In this work, we investigate the mirror diffusion of charged CR particles in a highly compressible environment using 3D-MHD simulations of an MC-like environment as the background and test particle propagation models to evaluate the diffusion behavior of the particles aiming at testing numerically the aforementioned predictions of LX21.
@@ 2.1 Background MHD turbulence
Figure 1 shows the power spectrum of the turbulent magnetic field for the background MHD model, after turbulence has reached a nearly steady state (see Barreto-Mota et al., 2021, for details). The blue line shows the result for the complete background magnetic field while the orange one shows the same model with a smoothed background magnetic field, where the smaller scale perturbations have been suppressed. To smooth the small scale perturbations, we calculated the 3D Fourier transform of each component of the magnetic field, set to zero the values corresponding to $k>5$, and then transformed it back to physical space. The resulting spectrum vanishes for $k>5$ (orange curve). From now on, we refer to this value as $k_{smooth}$. The green dashed line shows the Kolmogorov scaling with $k^{-5/3}$ for comparison. We see that a substantial part of the inertial range of the magnetic field in this smoothed model is eliminated.
In this study, we will employ both the smoothed and the complete background turbulent magnetic field models for the CR propagation analysis and compare their results. We note that smoothing the magnetic field significantly reduces the effects of resonant scattering on the propagation of particles with their gyroradii smaller than the smoothing scale, enabling us to highlight the effects of mirror diffusion. In the case when the particle Larmor radius $R_{L}$ is smaller than the damping scale of MHD turbulence, i.e., smoothed case, resonant scattering must be severely affected as the resonant condition is not satisfied, but non-resonant mirroring by larger-scale turbulent magnetic mirrors can still happen.
Figure 1: Magnetic power spectrum of the magnetic field components (solid lines). The blue line shows the result for the original complete magnetic field background (complete models as referred in the text). The orange line shows the same for the smoothed magnetic field background (see text for details). For reference, the vertical dotted lines show the corresponding value of k for each of the particle Larmor radius considered in this work.
@@ 2.2 Test Particle Simulations
To study the CR diffusion in the turbulent magnetic field generated in the aforementioned model, we employ the test particle code developed by Beresnyak et al. (2011) (see also Xu & Yan, 2013; Hu et al., 2022). The particles are allowed to diffuse through the statistically steady-state background turbulent magnetic field of the simulated model described in Section 2.1.
@@
where $R_{L}$ is the Larmor radius, $m$ the mass, $q$ the charge, and $\Omega$ the gyrofrequency. In this study, we consider the particle velocity given by the light speed $c$. We take particles with constant Larmor radius and initial $v_{\perp,0}\sim c$.
The code uses a Burlisch-Stoer method to trace the trajectories of test particles and uses an adaptive time step. At each time step, the magnetic field at the grid points is interpolated to the test particles position using a cubic spline routine [^2] The code uses $10^{3}$ neighboring points for the cubic spline interpolation. The large number of points allows for higher accuracy and smoother derivatives. (Press et al., 1986; Xu & Yan, 2013).
Given that the MHD turbulent models are scale-free and the energy of the particles depends on the properties of the simulated box (see Section 4), one can scale the system knowing the size of the domain $L_{0}$, the mean density and the sound speed as well as the scale independent parameters $M_{A}$ and $M_{s}$ that are defined at the injection scale of turbulence.
@@ 2.3 Initial setup
Each model is characterized by two free parameters, namely, the initial pitch angle ($\theta_{0}$) between the particles velocity and the local magnetic field, which we consider to be either $\mu_{0}=\cos\theta_{0}=0.2$ or $0.8$, and the Larmor radius ($R_{L}$), $R_{L}/L_{0}=0.03,0.06$ and $0.10$, in terms of the number of cells of the background simulation, $R_{L}(\#cells)=15,30$, and $50$, respectively (see the representation of these values in Figure 1). We note that particles are not accelerated and their energy stays the same in our simulations, we focus only on their diffusion. Each model name is labeled with these two parameters in Table 1. As described in section 2.1, we also consider two sets for the magnetic field configuration. One with the entire background turbulent magnetic field distribution, and another where the small-scale perturbations of the field were suppressed for $k_{smooth}=5$, which corresponds to a dissipative scale of about $\sim 0.2L_{0}$ ($\sim 100$ grid cells) [^3] Values of $k_{smooth}>5$ were tested, but the results showed little to no notable differences with regard to the model with the complete distribution of the magnetic field.. These latter models are labeled $k5$ in Table 1. The dissipation scale of the complete distribution of magnetic field is around $k\sim 250$, which corresponds to $0.004L_{0}$ approximately ($\lesssim 3$ grid cells). In all models, 3000 test particles are injected.
@@ 3.1 Particle trajectory and mirroring identification
Figure 2: Example of a test particle trajectory (represented by solid colored circles) with $R_{L}/L_{0}\simeq 0.03$ and $\mu_{0}=0.2$ in the background with the complete distribution of the magnetic field. Initially, the particle travels upwards parallel to the magnetic field (black circles), reentering the domain from the bottom face until it encounters magnetic mirrors that trap it for a while (dark purple to orange circles), diffusing perpendicularly to the field lines until it escapes again traveling downwards and reentering through the upper region of the domain, finding another mirror at the end of its trajectory (bright yellow circles). A few field lines are shown in red, white, and blue for visual distinction only. The density background is included to highlight the turbulent compressions of the medium. The color scales on the right hand side show the time elapsed in the particle trajectory between 850 and 1450 (same units as in Figure 3), and the density of the background (in code units), respectively. See Figure 3 and text for further details. We have plotted only a section of the cubic box in order to make the particle trajectory clearer.
@@
Figure 2 depicts an example of a particle trajectory in the original MHD turbulence simulation. A characteristic behavior observed in every model is long flights interrupted by successive bounces, where the particle interacts with different magnetic mirrors (Levy-flght-like propagation). As discussed in LX21, only particles with pitch angles above a critical value should interact with magnetic mirrors present along their path (Equation 4). When these conditions are met, the particle reverses its moving direction and bounces between mirrors. In the meantime, the particle also undergoes superdiffusion in the direction perpendicular to the mean magnetic field as a result of the turbulent reconnection of the magnetic fields. Because of the perpendicular superdiffusion, the particle does not remain trapped by mirrors, but diffusively move along the magnetic field when interacting with different mirrors [^4] As mentioned in Section 1, the effect of perpendicular superdiffusion has been thoroughly investigated elsewhere. We refer to numerical studies (e.g. Beresnyak, 2013; Xu & Yan, 2013; Eyink et al., 2013; Lazarian & Yan, 2014; Hu et al., 2022; Maiti et al., 2022; Zhang & Xu, 2023) that confirm the superdiffusive behavior of magnetic field lines and cosmic rays following the magnetic field (Lazarian & Vishniac, 1999; Lazarian & Yan, 2014). When the pitch angle becomes sufficiently small due to the pitch-angle scattering, the particle can escape from mirrors and undergo the scattering diffusion. With the stochastic change of pitch angle due to the gyroresonant scattering, the particle stochastically experiences both the mirror diffusion and scattering diffusion.
@@
Figure 3: Diagrams on the left (a) depict, for the complete background magnetic field distribution, the following quantities: top panel, the time evolution of the cosine of the pitch angle ($\mu$) for the same particle shown in Figure 2. The gradient color scheme depicts the portion of the particle trajectory as in Figure 2. Gray line shows the complete variation of $\mu$ over time and the blue line shows the same, but without the fastest variations of $\mu$. The middle panel shows the displacement of the particle parallel to the magnetic field with regard to its initial position, and the bottom panel shows $\zeta/\zeta_{avg}$ (Equation 3) over time. Red circles and black crosses are potential candidates for the beginning of bounces due to magnetic mirror interactions. Following the conservation of magnetic moment criteria introduced in the text, black crosses represent the rejected candidates. Time is normalized by the initial value of the girofrequency $\Omega=c/R_{L}$ (Equation 2), being $c=1$ in the code units. On the right side (b), the three first panels are the same as in (a) but zoomed in with a smaller range of $t$ marked by the vertical dashed lines. The bottom panel shows the magnetic field strength normalized by the mean magnetic field along the particle trajectory and the black circles indicate where we identify $\mu=0\,(\theta=90^{\circ})$ in the top panel (see text for details).
@@
In Figure 3, we show the time evolution of $\mu$ (top panel), the distance traveled parallel to the magnetic field versus time (middle panel), and $\zeta/\zeta_{avg}$ versus time, where $\zeta$ is the adiabatic invariant of the particle, which is conserved when the particle is bouncing in a mirror:
@@
and $\zeta_{avg}$ is the value of $\zeta$ calculated for $\mu=0.5$ and the average magnetic field $B_{0}$, for the same particle shown in Figure 2.
As an example, Figure 2 shows just a part of the particle’s trajectory. The respective time interval is shown in Figure 3 with the same color coding. Considering this time interval, the second panel of Figure 3 shows that the particle travels upward until it interacts with a mirror, staying more or less at the same height until it is able to continue its trajectory downwards after several consecutive mirroring events. In Figure 2 we see that during these several consecutive bounces, the particle continues to move perpendicular to the mean magnetic field, interacting with different mirrors. In the lower panel of Figure 3, the respective time interval also shows a nearly constant adiabatic invariant, which reinforces the fact that these are mirroring events.
The peaks of $\mu$ immediately before each bounce point (i.e., when $\mu=0$) in Figure 3 mark the beginning of a bounce, $\mu_{b}$, where $\mu_{b}$ is the value of $\mu$ at these peaks. This parameter represents the angle at which the particle begins interacting with a magnetic mirror. Specifically, if $|\mu_{b}|>|\mu_{c}|$, the particle should not be subject to mirroring,where $\mu_{c}=\cos{\theta_{pitch,c}}$ is the critical pitch-angle cosine threshold for mirror diffusion, as defined in Section 1. Unless otherwise noted (e.g., in Figures 3 and 4), all analyses use the absolute value of $\mu_{b}$.
@@
To identify the peaks in $\mu$ in Figure 3, we first filter the values of $\mu$ along time using a Butterworth digital filter (Virtanen et al., 2020), which helps us to follow the average behavior of $\mu$ along time. Note that there is a fast variation of $\mu$ and that not all bounce points correspond to $\zeta/\zeta_{avg}$ constant. Considering the smaller segment of the particle’s trajectory depicted in the (b) panels on the right side of Figure 3, where the bounces are better spaced and vary roughly between $-\mu$ and $+\mu$, we find that $\mu=0$ also corresponds to maximum magnetic field $B/B_{0}$ values (see bottom panel). However, if we observe the region identified with a red arrow in the (b) panels, there is a change of direction that does not correspond to $\mu=0$. Due to data filtering in the top panel, we only find $\mu=0$ where the green arrow points in the bottom panel. This misidentification occurs because filtering $\mu$ obscures some mirror interactions. The same issue arises in the region indicated by the purple arrow in the second (b) panel from top.
@@
This variation is reduced comparing the first panels of Figures 3 and 4. Figure 4 shows the same information presented in Figure 3, but for a particle propagating in the smoothed background distribution of the magnetic field. In Figure 3, top panels, the gray line (original data) has a component that partially hides the average behavior of the particle which is depicted by the blue line (filtered data). In Figure 4, since the magnetic field background is smoothed and scattering is already reduced, the peaks are more easily identifiable. Not only that, but the changes in the particle direction clearly correspond to maximum values of $B/B_{0}$ in Figure 4(b). In summary, in the smoothed case, identifying the bounces is much easier because scattering at small scales is not involved and also there are fewer mirror structures at smaller scales. Nevertheless, a similar trend is observed in both the complete and smoothed cases, with $B/B_{0}$ maximums generally corresponding to $\mu=0$, as expected for mirroring.